The first law of thermodynamics is usually written as

Q=W+?U.

You need to look at the signs carefully; some other disciplines (chemistry, for example) may use other signs, which implies other definitions of these quantities.

1-Which physical law underlies the first law of thermodynamics?

-conservation of matter

-increase in entropy

-conservation of energy

-conservation of work

2- The quantity Q is the __________ that is added to the system.

-heat

- work

-entropy

-internal energy

3-The quantity W in the first law refers to the __________.

  

-work done on the system

-work done by the system

-change of energy in the system

-change of volume in the system

The quantity ?U in the first law of thermodynamics is the change in the __________ of the system.

-entropy

-internal energy

-heat energy

-temperature
Answer #1
Concepts and reason

The concept used to solve this problem is first law of thermodynamics.

Find the physical law underlies in the first law of thermodynamics.

Then find the explanation for the term QQ in first law of thermodynamics.

And also find the explanation for the term WW in the first law of thermodynamics.

Finally explain the term ΔU\Delta U in the first law of thermodynamics.

Fundamentals

The first law of thermodynamics says,

“The heat energy added to the closed system is equal to change in internal energy of the system and the amount of work done by the system.

Mathematical expression for the first law of thermodynamics is,

Q=ΔU+WQ = \Delta U + W

Here, ΔU\Delta U is the change in the internal energy, QQ is the heat added to the system, and WW is the work done by the system.

(1)

Mathematical expression for the first law of thermodynamics is,

Q=ΔU+WQ = \Delta U + W

Heat added to the system is utilized to increase the internal energy of the system and work done by the system.

The physical law underlies the first law of thermodynamics is,

• Conservation of energy.

Conservation is explained by first law of thermodynamics and it also referred as law of energy conservation.

(2)

Heat is energy that is added to the system increases its internal energy and aids it to do work.

From the mathematical expression for the first law of thermodynamics,

Q=ΔU+WQ = \Delta U + W

The correct options for QQ is, Heat.

(3)

According to first law,

Q=W+ΔUQ = W + \Delta U

The positive sign before the work done denotes the work done by the system and negative sign before the work done denotes the work done by the system.

The quantity WW in the first law refer,

• Work done by the system.

(4)

The change in internal energy for a closed system is equal to the heat supplied, minus the amount of work done on its surrounding by the system.

ΔU=QW\Delta U = Q - W

The correct options for ΔU\Delta U is Internal energy.

In first law of thermodynamics the term ΔU\Delta U denotes the change in internal energy.

Ans: Part 1

The physical law underlies the first law of thermodynamics is conservation of energy.